Calculate the amount of heat radiated per second by a body of surface area 12 cm2 kept in thermal equilibrium in a room at temperature 20°C. The emissivity of the surface = 0.80 and σ = 6.0 × 10-8 W/m2-K4.

Grade:11

1 Answers

Aditi Chauhan
9 years ago
Sol. A = 12 cm^2 = 12 × 10^–4 m2 T = 20°C = 293 K e = 0.8 σ = 6 × 10^–8 w/m^2-k^4 Q/t = Ae σT^4 = 12 × 10^–4 0.8 × 6 × 10^–8 (293)^4 = 4.245 × 10^12 × 10^–13 = 0.4245 ≈ 0.42